OpenStudy (anonymous):

can you explain me how to simplify: (A.)log10^(x) (B.) e^(lnx) (C.) ln(e^(2)ln(e ln e)) (D.) e^(2ln A- (ln B)/2) *everything that is raised to a power is in parenthesis :)

OpenStudy (akshay_budhkar):

A log10^x=xlog10=x (using log a^b=bloga

OpenStudy (akshay_budhkar):

wait a sec i will get u reliable help lol

OpenStudy (anonymous):

thank you so much!

OpenStudy (akshay_budhkar):

hi joe

OpenStudy (anonymous):

>.>

OpenStudy (akshay_budhkar):

sorry for troubling.. just do this one

OpenStudy (anonymous):

\[e^{\ln x}= x, \ln e ^{x}, x>0\] by definition

OpenStudy (akshay_budhkar):

the answer for 2nd one is right.. e^lnx=x.... The property i told you was right too :)

OpenStudy (akshay_budhkar):

now using the first proprty the answer to the third one is 2

OpenStudy (akshay_budhkar):

i gotta leave urgent sorry bye

OpenStudy (anonymous):

kk

OpenStudy (akshay_budhkar):

sorry sparks were coming from the socket..